The scores in the summative test of 11-ABM A are normally distributed with a mean of 65 and a standard deviation of 12. Illustrate a normal curve and find the probability that some students got a score below 40. (Show your solution)



1
Expert's answer
2022-04-12T12:04:16-0400

Normal curve


"P(X<40)=P(N(65,12^2)<40)=P(65+12N(0,1)<40)=P(N(0,1)<-2.08)=0.01861"